Two containers contain two volume of same kind of ideal gas, the pressure and volume of one volume is  20.64 Pa, 2V0, the pressure 

and volume of another volume is  43.21 Pa and V0, the two container of gas are in thermal equalibrium with constant temperarture T0

external hot reserviour, the two containers are connected by a pipe which is initially blocked, when the block is removed and and 

two volume of gases mix, which of the following values is closest to the final equilibrium pressure? 


(A)   41.28

(B)   28.16

(C)   31.92

(D)   35.69

(E) All the above answers are incorrect



1
Expert's answer
2021-06-30T10:19:21-0400

"\mu_1=\frac{p_1V_1}{RT},"

"\mu_2=\frac{p_2V_2}{RT},"

"\mu=\frac{pV}{RT},"

"\mu=\mu_1+\mu_2,"

"p=\frac{p_1V_1+p_2V_2}{V_1+V_2},"

"p=\frac{20.64\cdot 2+43.21\cdot1}{2+1}=28.16~\text{Pa~(B)}."